OpenStudy (anonymous):

Algebra II problem Find the domain of (5x-(39)^(1/3))^1/2

OpenStudy (anonymous):

\[\sqrt{(5x-\sqrt[3]{39})}\]

OpenStudy (mertsj):

\[[5x-39^{\frac{1}{3}}]^{\frac{1}{2}}\]

OpenStudy (mertsj):

so 5x -39^1/3 cannot be negative. So 5x-39^1/3 must be greater than or equal to 0

OpenStudy (mertsj):

\[5x \ge 39^{\frac{1}{3}}\]

OpenStudy (anonymous):

the equations are coming out messed up now

OpenStudy (mertsj):